# Early-Exit and Instant Confidence Translation Quality Estimation

### v1

@@ Abstract
Quality estimation is omnipresent in machine translation, for both evaluation and generation. Unfortunately, quality estimation models are often opaque and computationally expensive, making them impractical to be part of large-scale pipelines. In this work, we tackle two connected challenges: (1) reducing the cost of quality estimation at scale, and (2) developing an inexpensive uncertainty estimation method for quality estimation. To address the latter, we introduce Instant Confidence COMET, an uncertainty-aware quality estimation model that matches the performance of previous approaches at a fraction of their costs. We extend this to Early-Exit COMET, a quality estimation model that can compute quality scores and associated confidences already at early model layers, allowing us to early-exit computations and reduce evaluation costs. We also apply our model to machine translation reranking. We combine Early-Exit COMET with an upper confidence bound bandit algorithm to find the best candidate from a large pool without having to run the full evaluation model on all candidates. In both cases (evaluation and reranking) our methods reduce the required compute by 50% with very little degradation in performance.
@@ 1 Introduction
Figure 1: Progression of predicted quality estimation score (dark line) and instant confidence estimation (shaded area) along the quality estimation model computation for four examples from the test set. Layer corresponds to compute cost. Red line stops computation because the confidence is high enough (early-exit).
@@
To overcome the first challenge, we propose Instant Confidence COMET, a model that jointly predicts translation quality and an associated uncertainty, without increasing the computational cost. We demonstrate in Section˜2 that the model can produce accurate confidence scores that strongly correlate with true prediction errors, without affecting the quality estimation performance.
We also use the confidence mechanism for our new Early-Exit COMET, a model that estimates translation quality already in the earlier layers. Examples of how score and confidence predictions develop throughout the layers are shown in Figure˜1. Early-Exit COMET uses the confidence score to determine whether an early prediction suffices or additional evaluation layers are needed. In Section˜3 we combine it with a simple threshold-based early-exit algorithm for faster sample-level quality estimation. Finally, in Section˜4 we use Early-Exit COMET for machine translation reranking. In this task, the machine translation model generates many candidates and the goal is to find the best one with as little compute as possible. We combine Early-Exit COMET with an upper confidence bound bandit algorithm for this reranking process and show that it outperforms strong random and logprob-based baselines.
@@ 2 Instant Confidence COMET
When a quality estimation model produces an output, such as 85, it is not clear whether this is just the model’s best guess or truly an accurate assessment of the quality. Having this additional information is important for decision making (Zhang et al., 2025; Farinhas et al., 2025), such as in commercial translation pipelines where high-scoring translations are marked as requiring very little human attention. If the high score is very uncertain, then it would be safer to have a human double-check these translations. Similarly, when designing budget-efficient human evaluation pipelines (Zouhar et al., 2025), human labor should be directed to examples where automated metrics are the least confident.
@@
A quality estimation model is a function $f$ that given a source $s$ and a translation $t$, computes an estimate of the quality: [^1] If a reference translation $r$ is also included in the input, then the quality score is called a reference-based metric.
@@ 2.1 Models
[…]
Notably, we do not backpropagate loss through $\hat{y}$ in the second term, only through $\hat{e}$. The architecture is illustrated in Figure˜2. During inference, the model has almost the same computational cost as an unmodified COMET model.
@@ 2.2 Results
[…]
@@
We compare our approach with previous work in Table˜1. The $\beta$ hyperparameter allows for trade-off between correlation with human score and confidence correlation with true error. For Monte Carlo dropout, as the number of runs is increased, the correlation between the model score variance (model confidence) and the true error increases. However, this comes at the very steep cost of having to run the same model with dropout multiple times, which is not feasible in many quality estimation applications. Moreover, using dropout during inference hurts the quality estimation performance, because part of the information flow in the network is obscured. Increasing the number of runs only partially compensates for this drop; indeed, even with 100 runs, the score correlation remains lower than that obtained with a single run with no dropout (0.281 vs. 0.327).
Upon first glance, the solution appears to be to run Monte Carlo dropout 100 times to obtain a good uncertainty estimate and once without the dropout to also obtain a good quality estimation. We show this approach as MC Dropout in Table˜1. However, due to the discrepancy between this model score and the uncertainty measure, the error correlation turns out to be very low, revealing this not to be a viable uncertainty prediction setup.
@@
Overall, our instant confidence approach is the best in terms of cost and competitive in terms of human and confidence correlations. In Appendix Figure˜7 we see that on average when the quality estimation model has high confidence (low $\hat{e}$) then it is likely to have an accurate prediction (low $|\hat{y}-y|$). Notably, the lowest confidence bin corresponds to samples where the quality estimation model is very incorrect ($|\hat{y}-y|>20$).
@@ 3.1 Early-Exit COMET with Self-Confidence
The architecture is illustrated in Figure˜3.
@@
Table 2: Pearson correlations between intermediate layer outputs (green) and between intermediate layer outputs and humans (purple) for supervised Early-Exit as described in Section˜3.1 (left) and unsupervised Early-Exit based on standard COMET (right). See Appendix Section˜3.1 for detailed version.
@@ 3.2 Results
We now discuss our findings with Early-Exit COMET. We use the same experimental setup as in Section˜2.
@@
To measure the quality of the early layer scores, we calculate correlations with the final layer scores, as well as with human evaluations. For comparison, we include the baseline COMET model, applying the final layer regressor to the intermediate embeddings to get intermediate scores. In Section˜3.1 we show that earlier layer scores of baseline COMET do not correlate strongly with final layer scores or with human judgments. However, with direct supervision at each layer, we see much better results for Early-Exit COMET. At layer 5 we already see a correlation score of $0.65$ with the final layer and $0.207$ with human scores. By layer $13$, the correlation with human scores is $0.278$, comparable to the final layer. We include a version of Early-Exit COMET with separate regression heads for each layer in Appendix Section˜3.1, but we do not observe any improvements.
To measure the quality of the self-confidence error predictions, we plot the average predicted error versus the true error in Appendix Figure˜6. We also include correlation scores for selected layers showing the correlation between the predicted and true errors, e.g., $0.44$ for layer $9$ scores. This enables early-exit decision making, which we introduce in the next section.
@@ 3.3 Deciding When to Early-Exit
In some cases, the Early-Exit COMET model is already close to the final assessment after a few layers (top row in Figure˜1). In these cases, we do not need to continue the computation if we are confident that the final outcome will be very close. The per-layer confidences can inform this decision.
In Algorithm˜1, we outline a simple heuristic that stops the Early-Exit COMET computation when the error prediction is low enough. To evaluate this algorithm, we compare it with two baselines that do not use the confidence scores: (1) Constant-Exit (Appendix, Algorithm˜3) stops at a constant, pre-defined layer; and (2) Variance-Exit (Appendix, Algorithm˜4) stops when the variance of three consecutive predictions is under a chosen threshold.
@@
3: Compute $L_{i}(s,t)$ from $L_{i-1}(s,t)$ $\triangleright$ next layer
[…]
@@
Results.
Figure˜4 shows the performance of Algorithm˜1 versus the baselines. We plot the correlation with final layer scores and human judgments for different budgets (total computation costs). The cost is relative to calculating the full COMET score for all inputs ($100\%$). We vary the algorithm thresholds to explore different budgets.
@@ Results.
[…]
@@
[…]
@@
11:  $\mathcal{C}^{\prime}\leftarrow\mathcal{C}^{\prime}\setminus\{c^{*}\}$ $\triangleright$ Remove $c^{*}$ if fully evaluated
@@
[…]
@@
Figure 5: Quality of the candidates returned by the Upper Confidence Bound bandit. Quality is measured in terms of the average final candidate score and the proportion to top-1 candidates selected. We plot these measures for various evaluation budgets. Cost (or budget) is given relative to calculating the full COMET scores for all candidates ($100\%$). See results in tabular form in Appendix Table˜4.
@@ 4.1 Upper Confidence Bound Bandit
In our context, the “arms” are the translation candidates $\mathcal{C}$, and pulling an arm corresponds to calculating an additional quality estimation model layer. Since the predictions improve with more layers (Section˜3.1), each “pull” of the “arm” improves our estimate of the reward for the given candidate. For each candidate, we always consider the reward of the highest layer explored thus far and its associated confidence score. The computation budget determines the number of pulls available. When the budget runs out, we pick the candidate with the highest reward estimate, ignoring the associated confidence scores. A more precise formulation is provided in Algorithm˜2. The $\gamma$ controls the balance between exploration and exploitation. We use $\gamma=1$ for most experiments, but an ablation study can be found in Appendix˜E.
@@ 4.2 Results
Section˜3.1 shows the quality of the bandit output as we increase the evaluation budget. We plot both the average score of the final candidates and the rate at which the top-1 (best full COMET score) candidate is selected.
Surprisingly, the logprob-based baselines underperform the random baseline in almost all settings, in both metrics, and for all budgets. We hypothesize that this is due to lower diversity in the subset of candidates selected. In contrast, the bandit outperforms the random baseline in almost all scenarios and for all budgets. In particular, with $60\%$ of the compute budget, there is almost no drop in the average translation quality. This indicates both that the Early-Exit COMET scores and error predictions are valuable and that the upper confidence bound bandit is able to make efficient use of these estimates.
For more detailed results on the upper confidence bound bandit, please see the ablation studies in Appendix˜E.
@@ 5 Related Work
Quantifying uncertainty in learned automated machine translation metrics was first proposed by Glushkova et al. (2021), who use the variance from ensembles or Monte Carlo dropout as a measure of uncertainty. A different approach by Zerva et al. (2022) introduces a secondary confidence estimation model to complement the original quality estimation model. [^6] They also propose an approach whereby one model predicts both quantities, similar to our instant confidence. See Appendix B documenting our attempts at replicating this model. Instead, we optimize a single model that produces two scores: a quality estimate and a measure of uncertainty. We optimize this model jointly in each training step, optionally optimizing for predictions at each layer and optionally with respect to the last layer’s prediction. Finally, we note that the focus of Zerva et al. (2022) is on analyzing the source of uncertainty, while our focus is on using uncertainty to make quality estimation more reliable and efficient.
@@
Multiple previous works investigated improving the efficiency of calculating trained metrics. For large scales, Rei et al. (2022a) use length-batching to speed up inference. At the same time, they statically prune the quality estimation model, which is similar to our constant-exit approach. Cheng et al. (2024) use a smaller baseline language model rather than the default XLM-Roberta for COMET. Zouhar et al. (2024) find that simply quantizing the model to half precision has almost no effect on the final quality estimation performance while halving the compute costs. Gowda et al. (2023, 2024) port COMET to a faster inference engine for massive speed gains. Finally, Larionov et al. (2024) explore pruning, distillation, and quantization for a very large quality estimation model, xCOMET (Guerreiro et al., 2024). All of these approaches are orthogonal to our method and could be used in combination.
@@
Previous works have proposed methods to stop computation at a particular Transformer block layer (Zhou et al., 2020), but these methods are largely applicable to classification tasks. Xin et al. (2021) propose learning-to-exit, which we loosely follow in our work. However, instead of predicting a probability of success, we predict the absolute error of the model, which is directly interpretable.

### v2

@@ Abstract
Quality estimation is omnipresent in machine translation, for both evaluation and generation. Unfortunately, quality estimation models are often opaque and computationally expensive, making them impractical to be part of large-scale pipelines. In this work, we tackle two connected challenges: (1) reducing the cost of quality estimation at scale, and (2) developing an inexpensive uncertainty estimation method for quality estimation. To address the latter, we introduce Instant Confidence COMET, an uncertainty-aware quality estimation model that matches the performance of previous approaches at a fraction of their costs. We extend this to Early-Exit COMET, a quality estimation model that can compute quality scores and associated confidences already at early model layers, allowing us to early-exit computations and reduce evaluation costs. We also apply our model to machine translation reranking. We combine Early-Exit COMET with an upper confidence bound bandit algorithm to find the best candidate from a large pool without having to run the full evaluation model on all candidates. In both cases (evaluation and reranking) our methods reduce the required compute by 50% with very little degradation in performance. Finally, we show how Instant Confidence COMET can be used to decide which translations a human evaluator should score rather than relying on the COMET score.
@@ 1 Introduction
Figure 1: Progression of predicted quality estimation score (dark line) and instant confidence estimation (shaded area) along the quality estimation model computation for four examples from the test set. Layer corresponds to compute cost. Red line | stops computation because the confidence is high enough (early-exit).
@@
To overcome the first challenge, we propose Instant Confidence COMET, a model that jointly predicts translation quality and an associated uncertainty, without increasing the computational cost. We demonstrate in Section 2 that the model can produce accurate confidence scores that strongly correlate with true prediction errors, without affecting the quality estimation performance.
We also use the confidence mechanism for our new Early-Exit COMET, a model that estimates translation quality already in the earlier layers. Examples of how score and confidence predictions develop throughout the layers are shown in Figure 1. Early-Exit COMET uses the confidence score to determine whether an early prediction suffices or additional evaluation layers are needed. In Section 3 we combine it with a simple threshold-based early-exit algorithm for faster sample-level quality estimation. Finally, in Section 4 we use Early-Exit COMET for machine translation reranking. In this task, the machine translation model generates many candidates and the goal is to find the best one with as little compute as possible. We combine Early-Exit COMET with an upper confidence bound bandit algorithm for this reranking process and show that it outperforms strong random and logprob-based baselines.
Another use for confidence scores by Instant Confidence COMET is to decide whether a particular translation needs to be scored by humans or if the automated metric suffices. We show positive results for this mixed human-metric annotation in Section 5.
@@ 2 Instant Confidence COMET
When a quality estimation model produces an output, such as 85, it is not clear whether this is just the model’s best guess or truly an accurate assessment of the quality. Having this additional information is important for decision making (Zhang et al., 2025; Farinhas et al., 2025), such as in commercial translation pipelines where high-scoring translations are marked as requiring very little human attention. If the high score is very uncertain, then it would be safer to have a human double-check these translations. Similarly, when designing budget-efficient human evaluation pipelines (Zouhar et al., 2025b), human labor should be directed to examples where automated metrics are the least confident.
@@
A quality estimation model is a function $f$ that given a source $s$ and a translation $t$, computes an estimate of the quality:111 If a reference translation $r$ is also included in the input, then the quality score is called a reference-based metric.
@@ 2.1 Models
$\mathcal{L}_{2}(y,\hat{y})+\beta\cdot\mathcal{L}_{2}(|y-\hat{y}|,\hat{e})$ (4)
Notably, we do not backpropagate loss through $\hat{y}$ in the second term, only through $\hat{e}$. The architecture is illustrated in Figure 2. During inference, the model has almost the same computational cost as an unmodified COMET model.
@@ 2.2 Results
$\bm{\downarrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$
Instant Confidence $\beta{=}0.25$ 1$\times$ 0.316 0.222
Instant Confidence $\beta{=}0.5$ 1$\times$ 0.309 0.224
Instant Confidence $\beta{=}0.75$ 1$\times$ 0.326 0.228 $\bigstar$
Instant Confidence $\beta{=}1.0$ 1$\times$ 0.330 0.207
Instant Confidence $\beta{=}1.5$ 1$\times$ 0.325 0.200
No Confidence 1$\times$ 0.327 -
MC Dropout (2) 2$\times$ 0.210 0.201
MC Dropout (5) 5$\times$ 0.247 0.267
MC Dropout (10) 10$\times$ 0.262 0.301
MC Dropout (50) 50$\times$ 0.279 0.328
MC Dropout (100) 101$\times$ 0.281 0.333
MC Dropout (2) 3$\times$ 0.327 0.061
MC Dropout (5) 6$\times$ 0.327 0.092
MC Dropout (10) 11$\times$ 0.327 0.115
MC Dropout (50) 51$\times$ 0.327 0.131
MC Dropout (100) 101$\times$ 0.327 0.134
DUP (parallelizable) 2$\times$ 0.327 0.135
DUP (sequential) 2$\times$ 0.327 0.216
@@
We compare our approach with previous work in Table 1. The $\beta$ hyperparameter allows for trade-off between correlation with human score and confidence correlation with true error. For Monte Carlo dropout, as the number of runs is increased, the correlation between the model score variance (model confidence) and the true error increases. However, this comes at the very steep cost of having to run the same model with dropout multiple times, which is not feasible in many quality estimation applications. Moreover, using dropout during inference hurts the quality estimation performance, because part of the information flow in the network is obscured. Increasing the number of runs only partially compensates for this drop; indeed, even with 100 runs, the score correlation remains lower than that obtained with a single run with no dropout (0.281 vs. 0.327).
Upon first glance, the solution appears to be to run Monte Carlo dropout 100 times to obtain a good uncertainty estimate and once without the dropout to also obtain a good quality estimation. We show this approach as MC Dropout in Table 1. However, due to the discrepancy between this model score and the uncertainty measure, the error correlation turns out to be very low, revealing this not to be a viable uncertainty prediction setup.
@@
Overall, our instant confidence approach is the best in terms of cost and competitive in terms of human and confidence correlations. In Appendix Figure 8 we see that on average when the quality estimation model has high confidence (low $\hat{e}$) then it is likely to have an accurate prediction (low $|\hat{y}-y|$). Notably, the lowest confidence bin corresponds to samples where the quality estimation model is very incorrect ($|\hat{y}-y|>20$).
@@ 3.1 Early-Exit COMET with Self-Confidence
The architecture is illustrated in Figure 3.
@@
Table 2: Pearson correlations between intermediate layer outputs (green) and between intermediate layer outputs and humans (pink) for supervised Early-Exit as described in Section 3.1 (left) and unsupervised Early-Exit based on standard COMET (right). See Appendix Section 3.1 for detailed version.
@@ 3.2 Results
We now discuss our findings with Early-Exit COMET. We use the same experimental setup as in Section 2.
@@
To measure the quality of the early layer scores, we calculate correlations with the final layer scores, as well as with human evaluations. For comparison, we include the baseline COMET model, applying the final layer regressor to the intermediate embeddings to get intermediate scores. In Section 3.1 we show that earlier layer scores of baseline COMET do not correlate strongly with final layer scores or with human judgments. However, with direct supervision at each layer, we see much better results for Early-Exit COMET. At layer 5 we already see a correlation score of $0.65$ with the final layer and $0.207$ with human scores. By layer $13$, the correlation with human scores is $0.278$, comparable to the final layer. We include a version of Early-Exit COMET with separate regression heads for each layer in Appendix Section 3.1, but we do not observe any improvements.
To measure the quality of the self-confidence error predictions, we plot the average predicted error versus the true error in Appendix Figure 7. We also include correlation scores for selected layers showing the correlation between the predicted and true errors, e.g., $0.44$ for layer $9$ scores. This enables early-exit decision making, which we introduce in the next section.
@@ 3.3 Deciding When to Early-Exit
In some cases, the Early-Exit COMET model is already close to the final assessment after a few layers (top row in Figure 1). In these cases, we do not need to continue the computation if we are confident that the final outcome will be very close. The per-layer confidences can inform this decision.
In Algorithm 1, we outline a simple heuristic that stops the Early-Exit COMET computation when the error prediction is low enough. To evaluate this algorithm, we compare it with two baselines that do not use the confidence scores: (1) Constant-Exit (Appendix, Algorithm 3) stops at a constant, pre-defined layer; and (2) Variance-Exit (Appendix, Algorithm 4) stops when the variance of three consecutive predictions is under a chosen threshold.
@@
3:   Compute $L_{i}(s,t)$ from $L_{i-1}(s,t)$triangleright next layer
4:   $\hat{y}_{i},\hat{e}_{i}\leftarrow R(L_{i})$triangleright apply regressor head
@@
1:$\sigma\leftarrow e\cdot\sqrt{\frac{\pi}{2}}$triangleright Rescale the absolute error estimates (MAE) to standard deviation estimates
2:$S_{\mathcal{C}}\leftarrow\{(\hat{y}_{1}(c),\sigma_{1}(c))|c\in\mathcal{C}\}$triangleright Calculate and cache first layer COMET for all candidates
3:$\hat{B}\leftarrow|\mathcal{C}|$triangleright Initialize running evaluation costs
4:$\mathcal{C}^{\prime}\leftarrow\mathcal{C};\quad l(c)\leftarrow 1\ \forall c\in\mathcal{C}^{\prime}$triangleright Initialize remaining candidates and explored layers
@@
6:   $\mathrm{UCB}(c)\leftarrow\hat{y}_{l(c)}(c)+\gamma\sigma_{l(c)}(c)\quad\forall c\in\mathcal{C}^{\prime}$triangleright Calculate UCB for all (remaining) candidates
7:   $c^{*}\leftarrow\operatorname*{arg\,max}_{c\in\mathcal{C}^{\prime}}\mathrm{UCB}(c)$triangleright Choose candidate with highest UCB
8:   $l(c^{*})\leftarrow l(c^{*})+1;\quad\hat{B}\leftarrow\hat{B}+1$
9:   $S_{\mathcal{C}}\leftarrow S_{\mathcal{C}}\cup(\hat{y}_{l(c^{*})}(c^{*}),\sigma_{l(c^{*})}(c^{*}))$triangleright Calculate next layer COMET score for $c^{*}$
@@
11:      $\mathcal{C}^{\prime}\leftarrow\mathcal{C}^{\prime}\setminus\{c^{*}\}$triangleright Remove $c^{*}$ if fully evaluated
@@
13:end $\triangleright$ Return best candidate based on the most advanced
@@
Results.
Figure 4 shows the performance of Algorithm 1 versus the baselines. We plot the correlation with final layer scores and human judgments for different budgets (total computation costs). The cost is relative to calculating the full COMET score for all inputs ($100\%$). We vary the algorithm thresholds to explore different budgets.
@@ Results.
Figure 5: Quality of the candidates returned by the Upper Confidence Bound bandit. Quality is measured in terms of the average final candidate score and the proportion to top-1 candidates selected. We plot these measures for various evaluation budgets. Cost (or budget) is given relative to calculating the full COMET scores for all candidates ($100\%$). See results in tabular form in Appendix Table 4.
@@ 4.1 Upper Confidence Bound Bandit
In our context, the “arms” are the translation candidates $\mathcal{C}$, and pulling an arm corresponds to calculating an additional quality estimation model layer. Since the predictions improve with more layers (Section 3.1), each “pull” of the “arm” improves our estimate of the reward for the given candidate. For each candidate, we always consider the reward of the highest layer explored thus far and its associated confidence score. The computation budget determines the number of pulls available. When the budget runs out, we pick the candidate with the highest reward estimate, ignoring the associated confidence scores. A more precise formulation is provided in Algorithm 2. The $\gamma$ controls the balance between exploration and exploitation. We use $\gamma=1$ for most experiments, but an ablation study can be found in Appendix E.
@@ 4.2 Results
Section 3.1 shows the quality of the bandit output as we increase the evaluation budget. We plot both the average score of the final candidates and the rate at which the top-1 (best full COMET score) candidate is selected.
Surprisingly, the logprob-based baselines underperform the random baseline in almost all settings, in both metrics, and for all budgets. We hypothesize that this is due to lower diversity in the selected candidates subset. In contrast, the bandit outperforms the random baseline in almost all scenarios and for all budgets. In particular, with $60\%$ of the compute budget, there is almost no drop in the translation quality. This indicates that (1) the Early-Exit COMET scores and error predictions are valuable and (2) the upper confidence bound bandit makes efficient use of these estimates.
For more detailed results on the upper confidence bound bandit, please see the ablation studies in Appendix E.
5 Deferring to Humans
In this section, we use the confidence scores from Instant Confidence COMET to decide whether we can rely directly on the COMET score, or whether a translation should instead be scored by a human annotator.
Methods.
As the simplest baseline, we consider randomly selecting examples for human annotation. For a stronger baseline, as used in Zouhar et al. (2025a), we select the examples with the lowest scores for human annotation. This can be done with existing metrics. We compare this with our approach of human-annotating examples with the lowest confidence scores (highest predicted error). Finally, for both the low-score and low-confidence approaches, we also consider oracle versions where we prioritize examples with low human scores and high absolute error respectively.
Setup.
We analyze deferral rates ranging between 0% and 100%. To measure the success of the combined human-metric annotation, we look at how the final ranking of the machine translation model changes degrades as we score more of the data with the automated metric. We quantify this by the system-level Spearman correlation between ranking based on the combined annotation approach and 100% human annotation. We macro-average results across all considered languages in the dataset.
Figure 6: System-level correlation (macro Spearman) of a combined human-metric annotation with part of segments being annotated by humans. At 0%, all segments are scored by the metric. At 100%, all segments are scored by humans.
Results.
The results are shown in Figure 6. Our method outperforms random selection, though is on par with deferring low-scoring examples. However, lower scores correlate (0.243 macro Pearson) with the number of words in the translation. The length of the translation affects the human annotation time (Zouhar et al., 2025a) and thus e.g. 20% deferral with this approach might use more annotation budget than 20% deferral with random selection. On the other hand, low confidence does not correlate meaningfully (-0.087 macro Pearson) with the translation length, thus making it a safer option.
Finally, we note that there is a large gap between using the predicted confidence scores and the oracle confidence. This means, that improvements in estimating the automated metric error can lead to further gains in deferral efficiency.
@@ 6 Related Work
Quantifying uncertainty in learned automated machine translation metrics was first proposed by Glushkova et al. (2021), who use the variance from ensembles or Monte Carlo dropout as a measure of uncertainty. A different approach by Zerva et al. (2022) introduces a secondary confidence estimation model to complement the original quality estimation model.666 They also propose an approach whereby one model predicts both quantities, similar to our instant confidence. See Appendix B documenting our attempts at replicating this model. Instead, we optimize a single model that produces two scores: a quality estimate and a measure of uncertainty. We optimize this model jointly in each training step, optionally optimizing for predictions at each layer and optionally with respect to the last layer’s prediction. Finally, we note that the focus of Zerva et al. (2022) is on analyzing the source of uncertainty, while our focus is on using uncertainty to make quality estimation more reliable and efficient. We direct the reader to the work of Lahlou et al. (2023) for more theoretical yet comprehensive treatment of directly estimating model confidence.
@@
Multiple previous works investigated improving the efficiency of calculating trained metrics. For large scales, Rei et al. (2022a) use length-batching to speed up inference. At the same time, they statically prune the quality estimation model, which is similar to our constant-exit approach. Cheng et al. (2024) use a smaller baseline language model rather than the default XLM-Roberta for COMET. Zouhar et al. (2024) find that simply quantizing the model to half precision has almost no effect on the final quality estimation performance while halving the compute costs. Gowda et al. (2023, 2024) port COMET to a faster inference engine for massive speed gains. Larionov et al. (2024) explore pruning, distillation, and quantization for a very large quality estimation model, xCOMET (Guerreiro et al., 2024). All of these approaches are orthogonal to our method and could be used in combination.
@@
Many works have explored intermediate model predictions (Liu et al., 2019; Belrose et al., 2023, inter alia). However, the key ingredient is to know when to stop the computation, such as at a particular Transformer block layer (Zhou et al., 2020), but these methods are largely applicable to classification tasks. Xin et al. (2021) propose learning-to-exit, which we loosely follow in our work. However, instead of predicting a probability of success, we predict the absolute error of the model, which is directly interpretable.